The electromagnetic wave that delivers a cellular phone call to a car has a magnetic field with an rms value of 1.8 × 10-10 T. The wave passes perpendicularly through an open window, the area of which is 0.23 m2. How much energy does this wave carry through the window during a 32-s phone call?
Answer #1

S = total energy / time*area

S = U / t*A

U = (1.8*10^-10)(0.23)(33)

= 1.324*10^-9 J